OpenStudy (jenniferjuice):

Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (campbell_st):

its as simple as \[x^2 + x + 1 = 0\] the general form of a quadratic is \[ax^2 + bx + c = 0\] match the coefficients

OpenStudy (jenniferjuice):

so a,b,and c would all be one?

OpenStudy (jenniferjuice):

like this? 1x^2+1x+1 a=1 b=1 c=1

OpenStudy (campbell_st):

yep thats it

OpenStudy (campbell_st):

the equation has complex solutions.... and the discriminant is used to determine the types of solutions \[\Delta = b^2 - 4ac\]

OpenStudy (jenniferjuice):

wait but would i solve the question using that?

OpenStudy (campbell_st):

do you know the general quadratic formula...? \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] you know a, b and c so substitute and solve

OpenStudy (jenniferjuice):

okay ,yes i know it

OpenStudy (jenniferjuice):

would i get x=1.581 ?